Bob, Tom and Phil are brothers. Altogether, their ages sum to 55. Tom is 4 years younger than Bob and Phil is 8 years older than Bob. How old is each brother?

Answer by ikleyn(52798)

.

Bob + Tom   + Phil    = 55


B   + (B-4) + (B + 8) = 55


3B + 4 = 55


3B     = 55 - 4


3B     = 51


 B     = 51/3 = 17.


ANSWER.  Bob is 17;  Tom is 17-4 = 13;  Phil is 17+8 = 25.


CHECK.   17 + 13 + 25 = 55.    ! Correct !

Solved.